OpenStudy (anonymous):

Solve 5x + 12 = x – 4.

OpenStudy (anonymous):

5x + 12 = x - 4 4x = -16 x = -4

OpenStudy (anonymous):

oh, I get it now!

OpenStudy (anonymous):

you get it? hope it helped, not really good at explaining